If the line passing through the points (a,2) and (7,7) is parallel to the line passing through the points (4,8) and (a+3,2) what is the value of a?
Answer by lwsshak3(11628)

If the line passing through the points (a,2) and (7,7) is parallel to the line passing through the points (4,8) and (a+3,2) what is the value of a?
***
parallel lines have equal slopes
slope=∆y/∆x
(7-2)/(7-a)=(2-8)/((a+3)-4)
5/(7-a)=-6/(a-1)
-42+6a=5a-5
a=37